# StructLM: Towards Building Generalist Models for Structured Knowledge Grounding

### v3

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the CodeLlama architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models (Xie et al., 2022) on 14 out of 18 evaluated datasets and establishes new SoTA performance on 7 SKG tasks. Furthermore, StructLM demonstrates strong generalization across 6 novel held-out SKG tasks, outperforming TableLlama by an average of 35% and Flan-UL2 20B by an avarage of 10%. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level. We release the model weights and training dataset to the community, along with relevant code on Github.
@@ 1 Introduction
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the CodeLlama family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin. Morever, we show that compared with TableLlama, a recent open language model finetuned for tabular inference tasks, our held-out performance is vastly superior. Compared with Flan-UL2-20B, our models also generalize better overall on the held out tasks by as much as 10% on average.
@@
  • We construct a large SKG instruction-tuning dataset with $1.1$ million samples. We train and release our 3 models that outperform the previous 3B USKG fine-tuned on individual tasks on a total of $14$ of $18$ tasks. StructLM also achieves SoTA results on $7$ of them.
@@ 4 Experiments
Dataset Metric SoTA ChatGPT Base ST FLAN-UL2 TableLlama USKG StructLM (Ours) $\Delta$
Size - - - 7B 7B×$\times$18 20B 7B 3B×$\times$18 7B 13B 34B -
@@
ToTTo BLEU 49.9 20.7 17.5 48.8 - - 49.0 49.4 49.3 50.2 +0.3
GrailQA EM 77.1 9.3 1.0 77.0 - - 70.1 80.4 79.2 82.2 +5.1
SQL2Text Blec 94.8 88.6 82.9 95.2 - - 94.8 93.8 88.5 92.6 -1.0
MMQA F1 85.3 59.6 30.7 81.5 - - 85.3 85.2 86.0 88.1 +2.8
Spider EM 80.5 43.8 5.2 67.3 - - 71.8 72.4 74.1 74.6 -5.9
KVRet All Micro 67.9 52.9 39.5 70.9 - - 67.9 72.6 69.5 69.3 +4.7
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 - 59.4 59.2 59.1 61.1 -7.3
SParC EM 68.2 32.2 3.2 62.3 - - 61.5 61.9 64.9 63.4 -3.3
CompWebQ Acc 76.8 48.9 3.1 75.6 75.9 - 73.3 78.3 80.4 81.9 -5.1
TabFact Acc 93.0 62.4 0.0 79.6 87.1 82.5 83.7 80.8 84.7 86.6 -6.4
WikiTQ All Ex 65.9 24.8 0.2 45.7 54.6 - 49.3 50.1 53.4 55.7 -10.2
WikiSQL All Ex 93.0 31.5 0.4 86.5 87.3 - 86.0 88.7 87.2 87.6 -5.4
FeTaQA BLEU 39.0 7.4 5.6 33.8 35.8 39.0 33.4 36.0 35.6 37.5 -1.5
FEVEROUS Acc 85.6 57.8 58.4 78.1 85.6 - 82.4 84.4 85.0 85.7 +0.1
MultiWOZ Joint Acc 60.6 8.9 0.0 53.0 - - 55.4 54.5 53.0 53.8 -6.1
DART BLEU 52.0 59.0 54.6 60.3 50.4 - 46.7 62.2 61.4 61.8 +10.2
Logic2Text Blec 95.3 78.5 59.1 89.5 - - 91.4 88.9 90.1 89.1 -5.2
MTOP EM 87.5 1.4 0.0 77.4 87.5 - 86.8 81.2 81.6 82.1 -5.4
Average 74.9 39.5 20.2 68.2 - - 69.3 71.1 71.3 72.6 -2.2
@@
BIRD Acc 36.6* 21.8 0.0 24.4* 1.0 0 0 22.3 22.8 24.7 +2.9
CoSQL EM 58.3* 33.7 0.2 52.4* 5.1 0 0 49.8 52.2 55.0 +21.3
SQA Acc 70.5* 18.7 2.3 60.4* 70.1* 0 0 49.7 36.1 44.2 +31
Infotabs Acc 75.6* 46.9 40.2 68.7* 70.3 16.2 0 55.3 58.1 61.8 -7.1
WikiTableText BLEU 33.7* 3.8 5.7 39.8* 19.4 3.4 0 8.3 9.3 8.8 -9.9
Finqa Acc 71.1* 31.4 1.7 79.7* 5.9 2.6 0 27.3 25.6 36.2 +4.8
Average 57.6* 26.1 8.4 54.2* 28.6* 3.7 0 35.5 34.0 38.4 +10.3
Table 2: The overall evaluation results of our model against other baselines. Cells with ”-” in the held-in part mean that the model did not train on this dataset, and results are not comparable. USKG models are overfit to the held-in dataset labels, and thus cannot generalize comparably. Cells in the held-out section with ”*” are held-in results. SoTA results are copied from the original papers for reference. ST refers to the single-task fine-tuning result of CodeLlama-Instruct-7B on each dataset. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B, with the same structure input format as StructLM. $\Delta$ refers to the difference between StructLM and the best known result. score denotes the state-of-the-art score on specific tasks. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
When compared to the 18 task-specific USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although the single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B (Tay et al., 2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Our results on held-in datasets (Tabfact and FeTaQA) are on par with TableLlama (Zhang et al., 2023), which is an LLM pre-trained on 2.6M table understanding tasks. We are vastly outperforming TableLlama on the held-out datasets.

### v4

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on Mistral and the CodeLlama model family, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models (Xie et al., 2022) on 16 out of 18 evaluated datasets and establishes new SoTA performance on 8 SKG tasks. Furthermore, StructLM demonstrates strong generalization across 6 novel held-out SKG tasks, outperforming TableLlama by an average of 35% and Flan-UL2 20B by an average of 10%. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level. We release the model weights and training dataset to the community, along with relevant code on Github.
@@ 1 Introduction
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on Mistral and the CodeLlama family of code foundation models. When compared to USKG, we find that our models surpass these single-task models on $16$ of $18$ tasks. StructLM achieves SoTA on $8$ of 18 evaluated tasks, beating ChatGPT by a huge margin. Morever, we show that compared with TableLlama, a recent open language model finetuned for tabular inference tasks, our held-out performance is vastly superior. Compared with Flan-UL2-20B, our models also generalize better overall on the held out tasks by as much as 10% on average.
@@
  • We construct a large SKG instruction-tuning dataset with $1.1$ million samples. We train and release our models that outperform the previous 3B USKG fine-tuned on individual tasks on a total of $16$ of $18$ tasks. StructLM also achieves SoTA results on $8$ of them.
@@ 4 Experiments
Dataset Metric SoTA ChatGPT Base-M Base ST FLAN-UL2 TableLlama USKG StructLM (Ours) $\Delta$
Size - - - 7B 7B 7B×$\times$18 20B 7B 3B×$\times$18 7B-M 7B 13B 34B -
@@
ToTTo BLEU 49.9 20.7 17.9 17.5 48.8 - - 49.0 49.8 49.4 49.3 50.2 +0.3
GrailQA EM 77.1 9.3 1.5 1.0 77.0 - - 70.1 81.2 80.4 79.2 82.2 +5.1
SQL2Text Blec 94.8 88.6 90.7 82.9 95.2 - - 94.8 95.2 93.8 88.5 92.6 +0.4
MMQA F1 85.3 59.6 41.5 30.7 81.5 - - 85.3 85.5 85.2 86.0 88.1 +2.8
Spider EM 80.5 43.8 31.0 5.2 67.3 - - 71.8 72.4 72.4 74.1 74.6 -5.9
KVRet All Micro 67.9 52.9 34.4 39.5 70.9 - - 67.9 72.2 72.6 69.5 69.3 +4.7
HybridQA Acc 68.4 23.7 12.9 2.3 58.4 61.0 - 59.4 62.6 59.2 59.1 61.1 -5.8
SParC EM 68.2 32.2 23.7 3.2 62.3 - - 61.5 63.3 61.9 64.9 63.4 -3.3
CompWebQ Acc 76.8 48.9 30.9 3.1 75.6 75.9 - 73.3 79.9 78.3 80.4 81.9 +5.1
TabFact Acc 93.0 62.4 25.7 0.0 79.6 87.1 82.5 83.7 84.6 80.8 84.7 86.6 -6.4
WikiTQ All Ex 65.9 24.8 6.7 0.2 45.7 54.6 - 49.3 56.8 50.1 53.4 55.7 -9.1
WikiSQL All Ex 93.0 31.5 21.5 0.4 86.5 87.3 - 86.0 87.0 88.7 87.2 87.6 -4.3
FeTaQA BLEU 39.0 7.4 13.7 5.6 33.8 35.8 39.0 33.4 37.5 36.0 35.6 37.5 -1.5
FEVEROUS Acc 85.6 57.8 73.2 58.4 78.1 85.6 - 82.4 85.9 84.4 85.0 85.7 +0.3
MultiWOZ Joint Acc 60.6 8.9 0.3 0.0 53.0 - - 55.4 55.4 54.5 53.0 53.8 -5.2
DART BLEU 52.0 59.0 47.4 54.6 60.3 50.4 - 46.7 63.2 62.2 61.4 61.8 +11.2
Logic2Text Blec 95.3 78.5 81.5 59.1 89.5 - - 91.4 89.5 88.9 90.1 89.1 -5.2
MTOP EM 87.5 1.4 0.8 0.0 77.4 87.5 - 86.8 75.8 81.2 81.6 82.1 -5.4
Average 74.9 39.5 30.8 20.2 68.2 - - 69.3 72.1 71.1 71.3 72.6 -1.2
@@
BIRD Acc 36.6* 21.8 11.5 0.0 24.4* 1.0 0 0 22.8 22.3 22.8 24.7 +2.9
CoSQL EM 58.3* 33.7 26.5 0.2 52.4* 5.1 0 0 52.8 49.8 52.2 55.0 +21.3
SQA Acc 70.5* 18.7 7.4 2.3 60.4* 70.1* 0 0 42.6 49.7 36.1 44.2 +31
Infotabs Acc 75.6* 46.9 49.1 40.2 68.7* 70.3 16.2 0 47.2 55.3 58.1 61.8 -8.5
WikiTableText BLEU 33.7* 3.8 3.9 5.7 39.8* 19.4 3.4 0 17.1 8.3 9.3 8.8 -2.3
Finqa Acc 71.1* 31.4 0.7 1.7 79.7* 5.9 2.6 0 29.5 27.3 25.6 36.2 +4.8
Average 57.6* 26.1 16.5 8.4 54.2* 28.6* 3.7 0 42.4 35.5 34.0 38.4 +8.2
Table 2: The overall evaluation results of our model against other baselines. 7B-M was trained with Mistral-7B as the base model. Cells with ”-” in the held-in part mean that the model did not train on this dataset, and results are not comparable. USKG models are overfit to the held-in dataset labels, and thus cannot generalize comparably. Cells in the held-out section with ”*” are held-in results. SoTA results are copied from the original papers for reference. ST refers to the single-task fine-tuning result of CodeLlama-Instruct-7B on each dataset. BASE and BASE-M refer to the 1-shot performance of CodeLlama-Instruct-7B, and Mistral-7B-Instruct-v0.2 respectively. $\Delta$ refers to the difference between StructLM and the best known result. score denotes the state-of-the-art score on specific tasks. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
When compared to the 18 task-specific USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B-M StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 16 of 18 datasets. It is worth noting that although the single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B (Tay et al., 2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Our results on held-in datasets (Tabfact and FeTaQA) are on par with TableLlama (Zhang et al., 2023), which is an LLM pre-trained on 2.6M table understanding tasks. We are vastly outperforming TableLlama on the held-out datasets.
@@
Additionally, our results in Table 2 support that the Mistral base model has stronger generalization and ability to handle SKG tasks than CodeLlama, with about a 10% advantage on metrics over all tasks. We can see that this strength transfers to the fine-tuned versions of the models.